OpenStudy (animalover02):

what is the slope of the line parallel to the line equation 2x-5y=9? A. -5/2 B. -2/5 C. 2/5 D. -9/5

OpenStudy (phi):

we have to change 2x-5y=9 into y= mx + b format

OpenStudy (animalover02):

i already thought C was the answer i was just haveing someone check my answer. is C correct?

OpenStudy (animalover02):

@phi

OpenStudy (animalover02):

im a freshman...but not smart in algebra

OpenStudy (will.h):

change 2x-5y=9 into slope intercept form by isolating y -5y = -2x + 9 y = (2/5)x + 9/-5 Since the lines are parallel that means the share the same slope slope above is 2/5

OpenStudy (animalover02):

Thank you so much!!!!!!!!!!!:)

OpenStudy (phi):

start with 2x-5y=9 add 5y to both sides 2x -5y + 5y = 5y+9 on the left side -5y + 5y add up to 0 we get 2x= 5y+9 now add -9 to both sides 2x - 9 = 5y +9-9 on the right side 9-9 is 0 we now have (switch the sides) 5y= 2x - 9 finally, divide both sides by 5 (that means both sides and *all terms*)

OpenStudy (joshmorris):

2x - 5y =9 5y = 2x -9 y = 2/5x -9/5 slope = 2/5 A. <---------------------(answer)

OpenStudy (will.h):

my pleasure

OpenStudy (animalover02):

;)